OpenStudy (anonymous):

Divide: (1 1/5) divided (14/25)

OpenStudy (anonymous):

mixed number divided by a fraction?

OpenStudy (anonymous):

Multiply (11/5)*(25/14)

OpenStudy (anonymous):

@completeidiot Yes, I believe so

OpenStudy (anonymous):

first convert the mixed number into an improper fraction

OpenStudy (anonymous):

@L-Lawliet-L No its divided, 1 and 1/5 divided by 14/25

OpenStudy (anonymous):

then use the fact that dividing by a fraction is the same as multiplying by its reciprocal as @L-Lawliet-L showed ex. \[\frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c}\]

OpenStudy (anonymous):

\[\frac{ 11 }{ 5 }\div \frac{ 14 }{ 25 }=\frac{ 11 }{ 5 }\times \frac{ 25 }{ 14 }\]

OpenStudy (anonymous):

its not eleven fifths, its one and one fifth

OpenStudy (anonymous):

@L-Lawliet-L Its one and one fifths

OpenStudy (anonymous):

anyways @taylorwristshot convert \[1 \frac{1}{5}\] into a improper fraction and what do you get?

OpenStudy (anonymous):

@completeidiot That would be 6/5, six fifths

OpenStudy (anonymous):

right so the problem is now \[\frac{6}{5} \div \frac{14}{25}\] now, dividing by a fraction is the same as multiplying by a reciprocal which means the above expression is the same as the expression below \[\frac{6}{5} \times \frac{25}{14}\]

OpenStudy (anonymous):

can you solve the expression now?

OpenStudy (anonymous):

@completeidiot Yes, wouldn't that make it 2 10/70?

OpenStudy (anonymous):

yea i would simplify the fraction though

OpenStudy (anonymous):

Yes, I did. Thank you! @completeidiot